alexgame:

A line passes through (4, 5) and (8, 9). Which equation best represents the line? y = x + 1 y = 1 y = 2x + 1 y = 5x + 8

mhanifa:

Your coordinates are (x, y) Substitute the values of x and y and see if you get a right equation. Another way is to make an equation using slope-intercept form: y = mx + b Find the slope first. Do you know the slope formula?

alexgame:

its going to be a

mhanifa:

@alexgame wrote:
its going to be a
5 = 4 + 1 9 = 8 + 1 Yes

snowflake0531:

Slope Formula\[m=\frac{y_{2} - y_{1}}{x_{2}-x_{1}}\] \[\frac{9-5}{8-4} = \frac{4}{4} = 1\] Slope=1 (4,5) \[5 = 4(1) +b \\ 5=4+b \\ b=1\] \[y=x+1\]